OpenStudy (anonymous):

Line AB contains points A (0, 1) and B (1, 5). The slope of line AB is (5 points) -4 4

OpenStudy (anonymous):

answers -4 -1/4 1/4 4

OpenStudy (anonymous):

i will fan and medal best answer

OpenStudy (trojanpoem):

5 - 1/ 1- 0 4/ 1 = 4

OpenStudy (texaschic101):

use slope formula : slope = (y2 - y1) / (x2 - x1) (0,1)....x1 = 0 and y1 = 1 (1,5)...x2 = 1 and y2 = 5 now sub and tell me what you get

OpenStudy (anonymous):

@texaschic101 i got 4 using slope fomula

OpenStudy (anonymous):